OpenStudy (anonymous):

what is the geometric mean of 8 and 18?

Nnesha (nnesha):

multiply the numbers together and then take square root of two numbers if there are 3 numbers then take 3rd root

Nnesha (nnesha):

\[\large\rm (product ~of ~all~ numbers )^\frac{ 1 }{ \color{reD}{number~of ~terms}~}\]

OpenStudy (anonymous):

k so if \[\sqrt{2^2X5}\it would be 2\sqrt{5}\]

Nnesha (nnesha):

is that different question?

Nnesha (nnesha):

a*

OpenStudy (anonymous):

not really, its random. i got the other one thx :)

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @WhiteFangIII what is the geometric mean of 8 and 18? \(\color{blue}{\text{End of Quote}}\) what did you get for this one ?

OpenStudy (anonymous):

12

Nnesha (nnesha):

okay right

OpenStudy (anonymous):

so bout the other ques, am i right?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

sweet thx